OpenStudy (anonymous):

If f(x)=3x+1 and f^-1=x-1/3 then f(2)

OpenStudy (anonymous):

f(x)=3x+1 f(y)=3y+1 f(a)=3a+1 i hope you see where this is going substitude the value in the parenthesis "( )" with x so f(2) = (3)(2)+1 or f(2) = 7

OpenStudy (anonymous):

ok I think I know want u did....thanks